OpenStudy (anonymous):

find the domain of the composite function f o g. f(x)=3/x-1 g(x)=2/x

OpenStudy (zzr0ck3r):

f composite g is f(2/x) = 3/(2/x) - 1 = (3/2)x-1 what is the domain of this?